OpenStudy (amonoconnor):

What is the first derivative (dy/dx) of the quantity that "y" equals, "2sinx - tanx"

OpenStudy (anonymous):

derivative of sin x is cos x derivative of tan x is sec^2(x)

OpenStudy (anonymous):

well, the derivative of 2 sin x will be 2 cos x So your final answer is 2 cos x - sec^2(x).

OpenStudy (anonymous):

You're welcome. @amonoconnor

OpenStudy (amonoconnor):

Thank you wonderful stranger! :)

OpenStudy (amonoconnor):

Oh wait.... When taking a derivative, wouldn't the 2 go away, because it's a constant... Not a term with x in it? So... Would it be : "cosxsex^2(x)" because... 2^(1-1)= = 2^(0) = (1) ...right?

zepdrix (zepdrix):

No silly! :) We ignore constant coefficients when we take a derivative.\[\Large\bf (2 \sin x)'\quad=\quad 2(\sin x)'\quad=\quad 2 (\cos x)\]But if you really wanted to convince yourself of that, yes you could take the derivative of the 2, but you'd have to do it using the product rule:\[\Large\bf (2\sin x)'\quad=\quad (2)'\sin x+2(\sin x)'\]\[\Large\bf =\quad (0) \sin x+2 (\cos x)\]

OpenStudy (amonoconnor):

You're freaking' amazing :) Thanks a bunch!